# Tel.: +91-080-2293 2058

### v1

@@ 2.2 Cross Attention Neural Operator Theory
We build the parametric map $\mathcal{G}_{\Theta}\colon\mathcal{A}\to\mathcal{H}$ as a composition of neural layers that perform lifting, iterative kernel integration, and projection as follows Li et al. (2023b).
$\mathcal{G}_{\theta}:=\mathcal{Q}\circ\sigma(\mathcal{W}_{T}+\mathcal{K}_{T}+b_{T})\circ\cdot\cdot\cdot\circ\sigma(\mathcal{W}_{1}+\mathcal{K}_{1}+b_{1})\circ\mathcal{P}\,,$ (4)
where $\mathcal{P}$ is the lifting operator, $\mathcal{Q}$ is the projection operator and the other terms are the $t=1,\cdots,T$ hidden layers that perform iterative kernel integration. Here, $\sigma$ is the pointwise nonlinear activation function, $\mathcal{W}_{t}$ is a local linear operator, $\mathcal{K}_{t}$ is the nonlinear kernel integral operator and $b_{t}$ is the bias function. These layers map each hidden representation to the next, i.e., $\{v_{t}\colon D_{t}\to\mathbb{R}^{d_{v_{t}}}\}\mapsto\{v_{t+1}\colon D_{t+1}\to\mathbb{R}^{d_{v_{t+1}}}\}$ using
@@
We propose an encoding layer $\mathcal{P}=W_{qpe}X$ for the lifting operator and a multi-layer perceptron $\mathcal{Q}=MLP(v_{T}(X))$ for the projection operator. For $\mathcal{K}_{t}$, we propose a cross-attention kernel integral operator
$\mathcal{K}_{t}=\int_{\partial\Omega}\frac{\text{exp}\left(\frac{\langle A_{h}v_{t}(X),B_{h}W_{bpe}X_{b}\rangle}{\sqrt{m}}\right)}{\int_{\partial\Omega}\text{exp}\left(\frac{\langle A_{h}v_{t}(X),B_{h}W_{bpe}X_{b}\rangle}{\sqrt{m}}\right)\rm{d}X_{b}}R_{h}W_{bve}b\,\rm{d}X_{b}\,,$ (6)
@@ 2.3 Practical Implementation
Fig. 1 illustrates our new Physics-Informed Transformer Neural Operator, named PINTO, that implements the cross-attention integral kernel integral operator. There are three stages of operation. In the first stage, we discretize the boundary function $\mathcal{B}(X_{b}^{i})$ at $L$ discrete points $X_{b}^{i}$, $i=1,\cdots,L$, and encode them using Boundary Value Encoding (BVE) and Boundary Position Encoding (BPE) units. We also encode the solution domain’s query point coordinate using Query Point Encoding (QPE) unit. All these encoding units are lifting operators as they encode scalar values as vectors of dimension $m$. We use dense layers in these units, but note that any other type of layer such as convolution or recurrent could also be used in the approach.
@@
$\zeta_{i}=\left(\sum_{l=1}^{L}\text{exp}\left(\frac{\langle A_{h}v_{t}(X)\;,\;B_{h}W_{bpe}(X_{b}^{l})\rangle}{\sqrt{m}}\right)\,\right)^{-1}\text{exp}\left(\frac{\langle A_{h}v_{t}(X),\,B_{h}W_{bpe}(X_{b}^{i})\rangle}{\sqrt{m}}\right)\,.$ (7)
@@
$v_{t+1}(X)=\sigma\left(A_{h}v_{t}(X)+\left(\sum_{h=1}^{H}\sum_{i=1}^{L}\zeta_{i}\cdot R_{h}W_{bve}(\mathcal{B}(X_{b}^{i}))\right)\right)\,.$ (8)
@@
During training, the data must be prepared such that the QPE unit is exposed to collocation points within the domain where the PDE loss is applied, as well as boundary points where both the boundary condition and the PDE loss are enforced. For each input to the QPE unit, the corresponding boundary conditions and the sequence of their values must be provided to the BPE and BVE units. In the case of Dirichlet boundary conditions, the boundary value is provided directly. However, for Neumann boundary conditions, the derivative values need to be fed to the BVE unit, and the loss function needs to be modified to account for these conditions. The code, PINTO architecture, and their trained weights have been released on GitHub with a Zenodo DOI Kumar (2024).
@@ 3 Applications of PINTO
The performance of PINTO in simulating the solution of IBVPs for unseen initial and boundary conditions is demonstrated by employing the following test cases: (i) $1$D-Advection equation, (ii) $1$D non-linear Burgers equation, (iii) $2$D unsteady Beltrami flow, (iv) $2$D steady Kovasznay flow, and (v) $2$D steady Lid-driven flow in the $x-y/z$ plane. The first two test cases demonstrate the applicability of our PINTO model in solving initial value problems and providing solutions for unseen times in the temporal domain. The latter ability is absent for the other leading physics-informed neural operators as we show. Numerical solutions of the advection equation from the PDEBENCH dataset (Takamoto et al., 2022), and those of the Burgers equation from an off-the-shelf solver Kovachki et al. (2021b) are used to validate the PINTO simulations. The next three test cases are governed by the Navier-Stokes equations. The third and fourth test cases, namely the Kovasznay and Beltrami flows, have analytical solutions (Kovasznay, 1948; Taylor, 1923) for different Reynolds numbers (Re), which are used to validate the PINTO solutions. These test cases demonstrate the applicability of our PINTO model to solve equations where the initial and boundary conditions vary spatially and temporally. In the fifth test case, called lid-driven cavity flow, we investigate the steady state flow within a vertical cavity where the top lid moves to the right at a constant speed. This flow is a simplified version of the wind-driven ocean flow and finds various practical applications in ocean and atmospheric studies. To validate our PINTO predictions, we obtain numerical solutions of the flow from a Finite Volume code (Ueckermann and Lermusiaux, 2012), where the boundary condition varied is the speed of the lid movement. Note that in all these situations, the validation data is not used for training. Training is performed using only Physics loss and initial/boundary conditions.
The performance of the PINTO model for all numerical examples is compared to the physics-informed DeepONet model (PI-DeepONet) (Lu et al., 2021a; Goswami et al., 2023). Although PI-DeepONet has not been used for learning generalized solutions of IBVPs, we repurposed them to find PDE solutions with different initial/boundary conditions by using the input functions for the branch net as the initial and boundary conditions. PI-DeepONet is chosen as the benchmark for comparison as it can be trained using only physics-loss unlike other neural operators that need some amount of simulation data. Table 1 summarizes the relative error for all test cases. In the upcoming sections, we take a deep dive into each test case. The code and data are available on Zenodo Kumar (2024) and Github.
@@
Beltrami Flow $0.53\%\,\left(0.1\%\right)$ $0.6\%\,\left(0.92\%\right)$ $2.62\%\,\left(4.19\%\right)$ $4.89\%\,\left(12.14\%\right)$
@@ 3.1 Advection Equation
$u(0,x)=\sum_{k_{i}=k_{1},\cdots,k_{N}}\left(A_{i}\sin{k_{i}x+\phi_{i}}\right),\,\text{in }\partial\Omega$ (9b)
where $x$ is the spatial coordinate, $t$ is the temporal coordinate, and $\beta$ is the constant advection speed. Different initial conditions are generated using the superposition of the sinusoidal wave equation (Eq. 9b), where $k_{i}=2\pi\{n_{i}\}/L_{x},i=1,\cdots,N$ whose wave numbers $\{n_{i}\}$ are random integers in $[1,n_{max}]$, $N$ is an integer that determines how many waves to add, $L_{x}$ is the domain size, $A_{i}$ is the amplitude randomly chosen in $[0,1]$, and $\phi_{i}$ is the phase randomly chosen in $(0,2\pi)$. We set $\beta=0.1$ for all experiments. We selected 100 initial conditions. For 80 of these initial conditions, the PINTO model is trained using the physics loss Eq. 3. The remaining 20 initial conditions are the unseen situations for which we will evaluate the initial condition generalization ability of PINTO. For validation, we use the PDEBENCH dataset that contains the numerical solutions of the advection equation for these initial conditions solved using a finite difference scheme with 2nd order upwind for time marching and a spatial grid size of $1024$ Takamoto et al. (2022). Note that the PINTO model is not trained using the PDEBENCH dataset.
@@
Figure 4: Learning curves during PINTO training for advection equation: a) Training (solid line) and validation (dashed line) loss curves for different learning rates. b) Training (solid line) and validation (dashed line) loss curves for different sequence lengths of initial conditions for a learning rate of $51e-$ 5.
Hyperparameter tuning was performed to select the number of CAUs, sequence length, learning rate, and activation functions using validation on unseen initial conditions, as described in Table 2. Figure 4 shows the learning curves of the PINTO model for various learning rates and sequence lengths. Figure 4 shows that lower learning rates and longer sequence lengths result in reduced validation loss. Consequently, the PINTO model was trained for $20000$ epochs using the Adam optimizer with a learning rate of $51e-$ 5 and a sequence length of $60$, achieving a mean relative error of $2.11\%$ on seen initial conditions and $2.85\%$ on unseen ones. For more details on the PINTO and PI-DeepONets hyperparameters, see A.
@@ 3.2 Burgers Equation
where $u_{0}$ is the initial conditions and $\nu$ is the viscosity coefficient. The initial conditions $u_{0}(x)\sim\mathcal{N}\left(0,\sigma^{2}\right)$ are sampled from a Gaussian random field with mean zero and covariance determined by the Laplacian Li et al. (2020a). We use periodic boundary conditions and set $\nu=0.01$. Here, PINTO is trained to learn the mapping between multiple initial conditions and the solution of the PDE. PINTO is trained for $80$ different initial conditions with physics loss (Eq. 3) considering $2000$ collocation, $250$ initial and boundary points in the domain $\left[0,1\right]\times\left[0,1\right]$ and tested on $20$ unseen initial conditions. Inference is performed on seen and unseen initial conditions for trained time and untrained time $t>1$. Numerical solutions for validation are obtained using an off the shelf solver Li et al. (2020a).
@@
Figure 5 shows the solution of the Burgers equation for two sets of seen and unseen initial conditions computed by PINTO. The corresponding numerical solution and the relative error are visualized to evaluate the performance of PINTO. Table 1 shows that the relative error for PI-DeepONet is three times that of our PINTO model for both seen and unseen initial conditions. Figure 6 shows the solution from PINTO, numerical solver, and PI-DeepONet at three discrete times $t=0.01,0.5,2$ to illustrate how the solution evolves in time. $t=2$ is a time step outside the temporal domain used to train all neural models. As time increases, the deviations of the solution computed by PINTO and PI-DeepONet from the numerical solution increase, but our PINTO model has fewer deviations compared to PI-DeepONet and has a better generalization for unseen initial conditions. The location with maximum value is marked with the letter A in Figure 5 to track how the solution evolves over time. We see that PINTO is able to maintain the extreme value of the solution better than PI-DeepONet. In addition, PINTO is able to maintain periodic boundary conditions better than PI-DeepONet.
A series of experiments similar to the Advection case have been performed to determine the number of CAUs for Burgers equations. Table. 2 summarizes the experiments with different learning rates and the number of CAUs, and its last column consists of the mean relative errors in the prediction of trained models on the set of unseen initial conditions.
@@ 3.3 Navier-Stokes Equation
u=g,on $\textbf{u}=\textbf{g},\,\text{on }\partial\Omega\,,$ (11c)
where $\mathbf{u}$ is the velocity vector, $p$ is the pressure, Re is the Reynolds number and $\mathbf{g}$ is the initial/boundary condition. For validation, we use the analytical solution of the Kovasznay and Beltrami flows, and the numerical solution from a finite volume solver for the lid driven cavity flow. In what follows, we discuss each test case in detail.
@@ 3.3.1 Kovasznay Flow
Figure 7: Kovasznay Flow: PINTO solutions (first row), corresponding analytical solution (second row) and relative error of PINTO solution (fourth row) for seen (first two columns) and unseen (last two columns) initial conditions (Re). Flow streamlines are overlaid on a background of the velocity magnitude $|V|=\sqrt{u^{2}+v^{2}}$ for the solution. Only magnitude of relative error is shown.
@@
$u(x,y)$ $=1-e^{\zeta x}\cos(2\pi y),$
$v(x,y)$ $=\frac{\zeta}{2\pi}e^{\zeta x}\sin(2\pi y),$
$p(x,y)$ $=\frac{1}{2}(1-e^{2\zeta x})\,,$
where $\zeta=0.5\text{Re}-\sqrt{0.25\text{(}Re)^{2}+4\pi^{2}}$. We consider the domain $\left[-0.5,1.0\right]\times\left[-0.5,1.5\right]$. To numerically solve the Navier-Stokes equations for the Kovasznay flow, boundary conditions are provided to the solver by evaluating the analytical solution at the boundary. Changing the value of Re changes the boundary condition. In this test case, PINTO learns the mapping between the boundary conditions and the solution of the PDE.

### v3

@@ 2.2 Cross Attention Neural Operator Theory
Consider the parametric map $\mathcal{G}_{\Theta}\colon\mathcal{A}\to\mathcal{H}$ as a composition of neural layers that perform lifting, iterative kernel integration, and projection as follows [25].
$\mathcal{G}_{\theta}:=\mathcal{Q}\circ\sigma(\mathcal{W}_{T}+\mathcal{K}_{T}+\gamma_{T})\circ\cdot\cdot\cdot\circ\sigma(\mathcal{W}_{1}+\mathcal{K}_{1}+\gamma_{1})\circ\mathcal{P}\,,$ (4)
where $\mathcal{P}$ is the lifting operator, $\mathcal{Q}$ is the projection operator and the other terms are the $t=1,\cdots,T$ hidden layers that perform iterative kernel integration. Here, $\sigma$ is the pointwise nonlinear activation function, $\mathcal{W}_{t}$ is a local linear operator, $\mathcal{K}_{t}$ is the nonlinear kernel integral operator and $\gamma_{t}$ is the bias function. These layers map each hidden representation to the next, i.e., $\{\mu_{t}\colon D_{t}\to\mathbb{R}^{d_{\mu_{t}}}\}\mapsto\{\mu_{t+1}\colon D_{t+1}\to\mathbb{R}^{d_{\mu_{t+1}}}\}$ using
@@
We propose an encoding layer $\mathcal{P}=MLP_{qpe}(X)$ for the lifting operator and a multi-layer perceptron $\mathcal{Q}=MLP(\mu_{T}(X))$ for the projection operator. For $\mathcal{K}_{t}$, we propose a cross-attention kernel integral operator
$\mathcal{K}_{t}=\int_{\partial\Omega}\frac{\text{exp}\left(\frac{\langle A_{h}\mu_{t}(X),B_{h}MLP_{bpe}\left(X_{b}\right)\rangle}{\sqrt{m}}\right)}{\int_{\partial\Omega}\text{exp}\left(\frac{\langle A_{h}\mu_{t}(X),B_{h}MLP_{bpe}\left(X_{b}\right)\rangle}{\sqrt{m}}\right)\rm{d}X_{b}}R_{h}MLP_{bve}\left(b\right)\,\rm{d}X_{b}\,,$ (6)
@@ 2.3 Practical Implementation
Fig. 1 illustrates our new Physics-Informed Neural Transformer Operator, named PINTO, that implements the cross-attention kernel integral operator. There are three stages of operation. In the first stage, we discretize the boundary function $\mathcal{B}(X_{b}^{i})$ at $L$ discrete points $X_{b}^{i}$, $i=1,\cdots,L$, and encode them using Boundary Value Encoding (BVE) and Boundary Position Encoding (BPE) units. We also encode the solution domain’s query point coordinate using Query Point Encoding (QPE) unit. All these encoding units are lifting operators as they encode scalar values as vectors of dimension $m$. We use dense layers in these units, but note that any other type of layer such as convolution or recurrent could also be used in the approach.
@@
$\zeta_{i}=\left(\sum_{l=1}^{L}\text{exp}\left(\frac{\langle A_{h}\mu_{t}(X)\;,\;B_{h}MLP_{bpe}(X_{b}^{l})\rangle}{\sqrt{m}}\right)\,\right)^{-1}\text{exp}\left(\frac{\langle A_{h}\mu_{t}(X),\,B_{h}MLP_{bpe}(X_{b}^{i})\rangle}{\sqrt{m}}\right)\,.$ (7)
@@
$\mu_{t+1}(X)=\sigma\left(A_{h}\mu_{t}(X)+\left(\sum_{h=1}^{H}\sum_{i=1}^{L}\zeta_{i}\cdot R_{h}MLP_{bve}(\mathcal{B}(X_{b}^{i}))\right)\right)\,.$ (8)
@@
During training, the data must be prepared such that the QPE unit is exposed to collocation points within the domain where the PDE loss is applied, as well as boundary points where both the boundary condition and the PDE loss are enforced. For each input to the QPE unit, the corresponding boundary conditions and the sequence of their values must be provided to the BPE and BVE units. In the case of Dirichlet boundary conditions, the boundary value is provided directly. However, for Neumann boundary conditions, the derivative values need to be fed to the BVE unit, and the loss function needs to be modified to account for these conditions. The code, PINTO architecture, and their trained weights have been released with a Zenodo DOI [78].
@@ 3 Applications of PINTO
The performance of PINTO in simulating the solution of PDEs for unseen initial and boundary conditions is demonstrated by employing the following test cases: (i) $1$D-Advection equation, (ii) $1$D non-linear Burgers equation, (iii) $2$D unsteady Beltrami flow, (iv) $2$D steady Kovasznay flow, and (v) $2$D steady Lid-driven flow in the $x-y/z$ plane. The first two test cases demonstrate the applicability of our PINTO model in solving initial value problems and providing solutions for unseen times in the temporal domain. The latter ability is absent for the other leading physics-informed neural operators as we show. Numerical solutions of the advection equation from the PDEBENCH dataset [79], and those of the Burgers equation from an off-the-shelf solver [80] are used to validate the PINTO simulations. The next three test cases are governed by the Navier-Stokes equations. The third and fourth test cases, namely the Kovasznay and Beltrami flows, have analytical solutions [81, 82] for different Reynolds numbers (Re), which are used to validate the PINTO solutions. These test cases demonstrate the applicability of our PINTO model to solve equations where the initial and boundary conditions vary spatially and temporally. In the fifth test case, called lid-driven cavity flow, we investigate the steady state flow within a vertical cavity where the top lid moves to the right at a constant speed. This flow is a simplified version of the wind-driven ocean flow and finds various practical applications in ocean and atmospheric studies. To validate our PINTO predictions, we obtain numerical solutions of the flow from a Finite Volume code [83], where the boundary condition varied is the speed of the lid movement. Note that in all these situations, the validation data is not used for training. Training is performed using only Physics loss and initial/boundary conditions.
The performance of the PINTO model for all numerical examples is compared to the physics-informed DeepONet model (PI-DeepONet) [7, 36]. Although PI-DeepONet has not been employed to learn generalized PDE solutions for unseen initial/boundary conditions akin to our problem, we adapted it to obtain PDE solutions under various initial/boundary conditions by treating the input functions for the branch network as initial and boundary conditions. PI-DeepONet is chosen as the benchmark for comparison as it can be trained using only physics-loss unlike other neural operators that need some amount of simulation data. Table 2 summarizes the relative error for all test cases. The definition of relative error and the results of other metrics to quantify the error are provided in  A. In the upcoming sections, we take a deep dive into each test case. The code and data for all test cases are available on Zenodo [78] and GitHub [^1] https://github.com/quest-lab-iisc/PINTO.
@@
Beltrami Flow $0.53\%\,\left(0.9\%\right)$ $0.6\%\,\left(0.92\%\right)$ $2.62\%\,\left(4.19\%\right)$ $4.89\%\,\left(12.14\%\right)$
@@ 3.1 Advection Equation
$u(0,x)=\sum_{\omega_{i}=\omega_{1},\cdots,\omega_{N}}\left(a_{i}\sin{\omega_{i}x+\phi_{i}}\right),\,\text{in }\partial\Omega$ (9b)
where $x$ is the spatial coordinate, $t$ is the temporal coordinate, and $\beta$ is the constant advection speed. Different initial conditions are generated using the superposition of the sinusoidal wave equation (Eq. 9b), where $\omega_{i}=2\pi\{n_{i}\}/\Delta_{x},i=1,\cdots,N$ whose wave numbers $\{n_{i}\}$ are random integers in $[1,n_{max}]$, $N$ is an integer that determines how many waves to add, $\Delta_{x}$ is the domain size, $a_{i}$ is the amplitude randomly chosen in $[0,1]$, and $\phi_{i}$ is the phase randomly chosen in $(0,2\pi)$. We set $\beta=0.1$ for all experiments. We selected 100 initial conditions. For 80 of these initial conditions, the PINTO model is trained using the physics loss Eq. 3. The remaining 20 initial conditions are the unseen situations for which we will evaluate the initial condition generalization ability of PINTO. For validation, we use the PDEBENCH dataset that contains the numerical solutions of the advection equation for these initial conditions solved using a finite difference scheme with 2nd order upwind for time marching and a spatial grid size of $1024$ [79]. Note that the PINTO model is not trained using the PDEBENCH dataset.
@@
Figure 4: Learning curves during PINTO training for advection equation: a) Training (solid line) and validation (dashed line) loss curves for different learning rates. b) Training (solid line) and validation (dashed line) loss curves for different sequence lengths of initial conditions (input to the BPE/BVE) for a learning rate of $51e-$ 5.
Hyperparameter tuning was performed to select the number of CAUs, sequence length, learning rate, and activation functions using validation on unseen initial conditions, as described in Table 5. Figure 4 shows the learning curves of the PINTO model for various learning rates and sequence lengths. Figure 4 shows that lower learning rates and longer sequence lengths result in reduced validation loss. This indicates that longer sequence lengths are generally better for achieving generalization to unseen initial/boundary conditions. However, lower learning rates require more epochs and longer sequences increase the compute time per epoch, leading to longer training time. Taking into account both factors, the PINTO model was trained for $20000$ epochs using the Adam optimizer with a learning rate of $51e-$ 5 and a sequence length of $60$, achieving a mean relative error of $2.11\%$ on the initial conditions seen and $2.85\%$ on the unseen conditions. For more details on the hyperparameters used for PINTO and PI-DeepONets, see B.
@@ 3.2 Burgers Equation
where $u_{0}$ is the initial conditions and $\nu$ is the viscosity coefficient. The initial conditions $u_{0}(x)\sim\mathrm{N}\left(0,\sigma^{2}\right)$ are sampled from a Gaussian random field with mean zero and covariance determined by the Laplacian [6]. We use periodic boundary conditions and set $\nu=0.01$. Here, PINTO is trained to learn the mapping between multiple initial conditions and the solution of the PDE. PINTO is trained for $80$ different initial conditions with physics loss (Eq. 3) considering $2000$ collocation, $250$ initial and boundary points in the domain $\left[0,1\right]\times\left[0,1\right]$ and tested on $20$ unseen initial conditions. Inference is performed on seen and unseen initial conditions for trained time and untrained time $t>1$. Numerical solutions for validation are obtained using an off the shelf solver [6].
@@
Figure 5 shows the solution of the Burgers equation for two sets of seen and unseen initial conditions computed by PINTO. The corresponding numerical solution and the relative error are visualized to evaluate the performance of PINTO. Table 2 shows that the relative error for PI-DeepONet is three times that of our PINTO model for both seen and unseen initial conditions. Figure 6 shows the solution from PINTO, numerical solver, and PI-DeepONet at three discrete times $t=0.01,0.5,2$ to illustrate how the solution evolves in time. $t=2$ is a time step outside the temporal domain used to train all neural models. As time increases, the deviations of the solution computed by PINTO and PI-DeepONet from the numerical solution increase, but our PINTO model has better predictions compared to PI-DeepONet and has a better generalization for unseen initial conditions. The location with maximum value is marked with the letter A in Figure 5 to track how the solution evolves over time. We see that PINTO is able to maintain the extreme value of the solution better than PI-DeepONet. In addition, PINTO is able to maintain periodic boundary conditions better than PI-DeepONet.
A series of experiments similar to the Advection case have been performed to determine the number of CAUs for Burgers equations. Table 5 summarizes the experiments with different learning rates and the number of CAUs, and its last column consists of the mean relative errors in the prediction of trained models on the set of unseen initial conditions.
@@
Figure 7 shows the performance of PINTO for $20$ different unseen initial conditions for longer times. Here, the mean relative error is computed for all points in the discrete spatial domain for a given time. It is observed that $62<t<$ 6 there is a rapid growth in the relative error, and beyond the time step $t>6$ all test cases are clustered into two groups; in one, the relative error is reaching saturation and in another, the relative error is still increasing, but with a slower rate. In all cases, the relative error is confined to 2-12%.
Figure 7: Burgers Equation: Evolution of mean relative error in time for $20$ different unseen initial conditions.
@@ 3.3 Navier-Stokes Equation
u=b,on $\textbf{u}=\textbf{b},\,\text{on }\partial\Omega\,,$ (11c)
where $\mathbf{u}$ is the velocity vector, $p$ is the pressure, Re is the Reynolds number and $\mathbf{b}$ is the initial/boundary condition. For validation, we use the analytical solution of the Kovasznay and Beltrami flows, and the numerical solution from a finite volume solver for the lid driven cavity flow. In what follows, we discuss each test case in detail.
@@ 3.3.1 Kovasznay Flow
Figure 8: Kovasznay Flow: PINTO solutions (first row), corresponding analytical solution (second row) and relative error of PINTO solution (fourth row) for seen (first two columns) and unseen (last two columns) initial conditions (Re). Flow streamlines are overlaid on a background of the velocity magnitude $|V|=\sqrt{u^{2}+v^{2}}$ for the solution. As the error is small, $\log_{10}$ of magnitude of the relative error is shown.
@@
$u(x,y)$ $=1-e^{\eta x}\cos(2\pi y),$
$v(x,y)$ $=\frac{\eta}{2\pi}e^{\eta x}\sin(2\pi y),$
$p(x,y)$ $=\frac{1}{2}(1-e^{2\eta x})\,,$
where $\eta=0.5\text{Re}-\sqrt{0.25\text{Re}^{2}+4\pi^{2}}$. We consider the domain $\left[-0.5,1.0\right]\times\left[-0.5,1.5\right]$. To numerically solve the Navier-Stokes equations for the Kovasznay flow, boundary conditions are provided to the solver by evaluating the analytical solution at the boundary. Changing the value of Re changes the boundary condition. In this test case, PINTO learns the mapping between the boundary conditions and the solution of the PDE.